An Unknown compound contains only 2 elements, platinum (Pt) and chlorine (Cl). S pure 5 g sample of the compound is analyzed and found to contain 2.105 g Cl.
A.) What percentage of the compound is Cl?
B.) Determine the empirical formula of the compound
C.) Gravimetric analysis confirms that in the 5 g sample there are .0148 mol present. Determine the molar mass of the compound.
D.) What is the molecular formula of this compound?
1
Expert's answer
2020-02-28T12:13:49-0500

А.

2.105 x 100/5 = 42.1%.

B. 2.105/35.5 : 2.895/195 = 4:1

PtCl4

C. 5/0.0148= 337.837838 g/mol.

D. PtCl4